This is not text book - statistics homework and it's rough!
The score on the entrance test for a well known law school has a mean score of 200 points and a standard deviation of 50 points. At value should the lowest passing score be set if the school wishes only 2.5% of those taking the entrance test to pass?

Answer by stanbon(48569)

The score on the entrance test for a well known law school has a mean score of 200 points and a standard deviation of 50 points.
What value should the lowest passing score be set if the school wishes only 2.5% of those taking the entrance test to pass?
----------------
Find the z-score associated with the highest 2.5% of the population.
z = 1.96
Find the x-value associated with that z-score.
z = (x-mu)/sigma
1.96 = (x-200)/50
x = 1.96*50 + 200
x = 298
-------------
That score separates the top 2.5% of the scores from the rest of
the population.
===============
Cheers,
Stan H.